OpenStudy (anonymous):

Two pizzas and a salad cost $29.00. Three pizzas and 2 salads cost $46.00. How much is a pizza? How much is a salad? Write the system of equations that would be used to solve this problem. none of the above

OpenStudy (anonymous):

two pizzas and a salad cost $29.00. Three pizzas and 2 salads cost $46.00. How much is a pizza? How much is a salad? Write the system of equations that would be used to solve this problem. Let x = price of a pizza y= price of a salad 2x +y=29.00 3x+2y=46.00 Solve the system of equations in problem #5. -4x-2y =-58.00 (Multiply the first equaton by -2.) 3x +2y =46.00 ---------------------------------- -x =-12 x=12.00 price of a pizza. From this equation: 2x+y=29.00 2(12)+ y=29.00 24+y=29.00 y=29=24 y=5 What is the cost of one salad? $4 $5 $6 $7 ANSWER: Cost of a salad is $5. What is the cost of one pizza in problem #5? $9 $10 $11 $12 ANSWER: Pizza costs $12.